# Gyroscopically stabilized quantum spin rotors

### v1

@@ Abstract
The free rotation of a magnetized body can differ significantly from that of the same object if it were non-magnetic. This, perhaps surprising, discrepancy is due to the fact that magnetic moments carry a spin angular momentum that contributes to the total angular momentum of the magnetized body, as quantified by the material-specific gyromagnetic ratio. Remarkably, it is the total angular momentum, i.e. the sum of mechanical and spin angular momentum, that enters the equations of motion of the body [1]. This gives rise to the so-called spin-rotation coupling [2] between rigid-body rotations and magnetic degrees of freedom, as famously witnessed in the Einstein-de Haas and Barnett effects [3, 4, 5, 6, 7]. The impact of spin-rotation coupling typically decreases with particle size, while it is still relevant for microscale and nanoscale magnetic objects [8, 9, 10, 11, 12, 13].
@@
Nanoscale particles of various shapes and with sizes ranging from hundreds of nanometers to tens of microns can be stably suspended in ultra-high vacuum using optical, electrical, or magnetic traps [14]. Monitoring and controlling their linear and rotational motion with laser fields enables precision sensing of forces and torques as required for fundamental tests of physics [15, 16]. Recent demonstrations of cooling their center-of-mass oscillations [17, 18, 19, 20] and their angular librations [21, 22] in an optical tweezer into the quantum regime constitute the first steps towards quantum applications of levitated particles. The demonstrated degree of control and read-out renders these systems ideal for performing quantum experiments with objects of unrivaled masses [23, 24, 25, 26, 27, 28].
@@
Figure 1: (a) A charged spheroidal nanodiamond with an embedded nitrogen-vacancy (NV) center is levitated in a Paul trap and revolves rapidly with frequency $\omega$ in the symmetry plane of the trap. The major principal axis $\mathbf{n}_{1}$ of the body is chosen to be aligned with the NV quantization axis so that the NV spin states $|S_{1}=\pm\hbar\rangle$ are split by the rotation-induced Barnett field as well as by an applied magnetic field $\mathbf{B}$. The particle rotation induces (b) an effective gyroscopic potential $I\omega^{2}/2\sin^{2}\beta$ stabilizing the out-of-plane rotation angle $\beta$, up to small librations $\xi$, while the external magnetic field together with spin-rotation coupling induce (c) a spin state-dependent potential $V_{m}(\gamma)$ for $\gamma$ rotations around the body-fixed symmetry axis depending on the spin states $|S_{1}=\hbar m\rangle$ with $m\in\{-1,0,1\}$.
@@
Experiments with levitated magnetic objects aim at harnessing their magnetic properties to gain better control over the internal and external degrees of freedom of the nanoparticle [29]. Magnetized particles have been trapped in electric [30, 31], magnetic [32, 33, 34], optical [35, 36], and superconducting traps [37, 38], where their center-of-mass and rotational motion can be monitored and manipulated. Various schemes have been devised to control their mechanical motion via driving the internal magnetization field [39, 40, 41, 42, 43] or individual magnetic defect centers, such as nitrogen-vacancy (NV) centers [44, 45, 29, 46, 47]. These schemes promise to enable sensing of magnetic fields and accelerations [48, 9] as well as the creation of massive superpositions [49, 26, 50, 51], and perhaps even probing the quantumness of gravity [52].
@@
In this article, we show how the spin-rotation coupling between a levitated nanoparticle and a single embedded spin degree of freedom can be observed and exploited for quantum experiments. The proposed scheme is based on the recent experimental demonstration of rapid spinning of electrically levitated nanodiamonds with embedded spin degrees of freedom [53, 31]. This spinning boosts the spin-rotation coupling while it also gyroscopically stabilizes the nanoparticle orientation in the co-rotating frame (see Fig. 1), so that a combination of magneto-static fields and microwave pulses then enables controlling the collective motion of hundreds of millions of atoms via a single spin.
@@
Specifically, we consider a symmetric nanodiamond with a single embedded NV spin levitated in the presence of an external static magnetic field $\mathbf{B}=B\mathbf{e}_{z}$, as schematically depicted in Fig. 1. The nanodiamond is approximated as an ellipsoidal rigid rotor with semi-axis lengths $l_{1}=l_{2}<l_{3}$, and moments of inertia $I_{1}=I_{2}=I$ with $I_{3}<I$. The nanoparticle orientation $\Omega=(\alpha,\beta,\gamma)$, parametrized by Euler angles in the $z$-$y^{\prime}$-$z^{\prime\prime}$ convention, specifies the tensor ${\rm R}(\Omega)$ which rotates from the lab-frame axis ${\bf e}_{j}$, $j=x,y,z$, to the rotor principal axes $\mathbf{n}_{\mu}(\Omega)={\rm R}(\Omega){\bf e}_{j}$, $\mu=1,2,3$. For instance ${\bf n}_{1}(\Omega)={\rm R}(\Omega){\bf e}_{x}$, etc.; an explicit expression for ${\bf n}_{\mu}(\Omega)$ is given in App. A. The inertia tensor depends on the orientation of the particle through $\mathrm{I}(\Omega)=\sum_{\mu}I_{\mu}\mathbf{n}_{\mu}(\Omega)\otimes\mathbf{n}_{\mu}(\Omega)$. A single NV center with spin-one angular momentum vector operator $\mathbf{S}$ is embedded in the nanodiamond, effectively introducing an internal spin triplet and adding the magnetic dipole moment $\gamma_{0}{\bf S}$ to the particle (with gyromagnetic ratio $\gamma_{0}>0$). The spin angular momentum adds to the mechanical angular momentum to yield the total angular momentum ${\bf J}$ of the system. As the spin quantization axis we choose $\mathbf{n}_{1}(\Omega)$, effectively breaking the symmetry of the particle shape; the NV zero-field splitting is $D_{\rm nv}\simeq 2\pi\times 2.87$ GHz.
@@
Gyroscopic stabilization is achieved by trapping a charged particle in an azimuthally symmetric quadrupole trap and spinning it with high rotation frequency $\omega$ around the lab-frame ${\bf e}_{z}$-axis, as recently realized in two experiments [53, 31]. For homogeneously charged particles the electric quadrupole tensor aligns with the inertia tensor, with quadrupole moments $Q_{1}=Q_{2}=Q$ and $Q_{3}>Q$. For vanishing electric dipole moment, the center-of-mass and rotational motion of the particle decouple so that the resulting rotational macromotion potential is determined by the ac voltage $U_{\rm ac}$, the electrode separation $d_{0}$, and the ac frequency $\omega_{\rm ac}$ [54]. The spin-rotor Hamiltonian is
@@
Here, $J_{\mu}={\bf J}\cdot{\bf n}_{\mu}(\Omega)$ denotes the body-frame component of the total angular momentum vector and likewise for ${\bf S}$. The confining quadrupole potential depends only on the polar angle $\beta$, implying that the lab-frame angular momentum component $p_{\alpha}={\bf J}\cdot{\bf e}_{z}$ is conserved. It is thus convenient to transform the spin-rotor state $|\psi\rangle$ to a co-rotating frame of boosted angular momentum via
@@
This adds the gyroscopic potential $I\omega^{2}/2\sin^{2}\beta$ [see Fig. 1(b) and App. B] to the quadrupole potential in Eq. (Gyroscopically stabilized quantum spin rotors), effectively confining the particle to the rotation plane $\beta\simeq\pi/2$ with harmonic frequency
@@
For large rotation rates $\omega_{\xi}\simeq\omega$, librations out of the rotation plane can be first linearized in $\xi=\pi/2-\beta$ and then adiabatically eliminated. Specifically, linearization yields
@@
with $p_{\gamma}=J_{3}$, $H_{\xi}=p_{\xi}^{2}/2I+I\omega_{\xi}^{2}\xi^{2}/2$ and we neglected non-dominant contributions to spin-rotation coupling as well as dispersion in the gyroscopic motion (see App. B). The fast librations out of the rotation plane can now be eliminated by neglecting transitions between eigenstates $|n\rangle$ of $H_{\xi}$ in a Born-Oppenheimer approximation (see App. C). This is achieved by expanding
@@
Figure 2: (a) Steady-state alignment $\langle\cos\gamma\rangle$ as a function of the applied magnetic field. The variance is indicated by the shaded region. (b) Potential energy surfaces $\Omega_{+}$ (solid) and $\Omega_{-}$ (dashed) for $B=-0.5\,\mathrm{mT}$ and $\omega/2\pi=1\,\mathrm{MHz}$. The inset shows the avoided crossing around $\gamma=\pi/2$ where the spin-rotor state $|\chi_{+}\rangle$ can be stably trapped in the $|\sigma_{x}=+1\rangle$ spin state. (c) Spin transition probability for an initial state in spin $|\sigma_{x}=+1\rangle$ (solid) and spin $|\sigma_{x}=-1\rangle$ (dashed) as a function of $\omega_{\gamma}t$ for different magnetic field strengths. The nanodiamond has an ellipsoidal shape with semi-axes $l_{3}=200$ nm and $l_{1}=l_{2}=0.3l_{3}$.
@@
$+\frac{1}{2}(\omega-\gamma_{0}B)\left[f_{n}(S_{1}-iS_{2})e^{-i\gamma}+{\rm h.c.}\right],$
@@
This Hamiltonian describes the coupling between the NV spin and $\gamma$-rotations with effective inertia $I_{\rm eff}=II_{3}/(I-I_{3})$. The final two terms are the adiabatic potential (or potential energy surface) for fixed libration state $|n\rangle$, containing the magnetic potential as well as spin-rotation coupling and the Barnett effect on the NV zero-field splitting. The adiabatic terms are weighted by the overlap integrals
$f_{n}=$ $\langle n|\exp\left(-\frac{ip_{\xi}}{\hbar}\frac{\hbar(\omega-\gamma_{0}B)}{I\omega^{2}}\right)|n\rangle$ (7a)
$g_{n}=$ $\langle n|\exp\left(-\frac{ip_{\xi}}{\hbar}\frac{2\hbar\gamma_{0}B}{I\omega^{2}}\right)|n\rangle$ (7b)
for transitions between displaced and non-displaced libration states. Note that $f_{n}\simeq 1$ for $n\ll I^{2}\omega^{4}/\hbar^{2}\gamma_{0}^{2}B^{2}$ and likewise $g_{n}\simeq 1$. Given that $\hbar\gamma_{0}B/I\omega^{2}\ll 1$ for realistic magnetic fields and rotation rates, the Hamiltonian (Gyroscopically stabilized quantum spin rotors) becomes approximately independent of $n$. Thus, all adiabatic spin-rotor states $|\chi_{n}\rangle$ evolve according to the same effective Hamiltonian
@@
This Hamiltonian is the main result of this work and the starting point for the proposals discussed below. It accounts for spin-rotation coupling in three different ways: First, the planar $\gamma$-rotation occurs with the effective moment of inertia $I_{\rm eff}$ because $\xi$-librations are gyroscopically stabilized and adiabatically eliminated. Second, the rotation rate $\omega$ of the particle induces the effective spin-dependent potential in the second line of the Hamiltonian (Gyroscopically stabilized quantum spin rotors). Third, the $\mathbf{B}$ field induces a coupling of the spin to rotations around the symmetry axis of the particle as described by the last term in the first line. Note that this coupling is typically small and will be neglected in the following. Next, we will study the implications of gyroscopic stabilization for the experimental observation of strong spin-rotation coupling.
(a) Barnett splitting— The Hamiltonian shows how the degeneracy of the $|S_{1}=\pm\hbar\rangle$ states is lifted by the Barnett effect for $B=0$. Specifically, neglecting $\gamma$-rotations by setting $\gamma=0$ and $p_{\gamma}=0$, the $|S_{1}=\pm\hbar\rangle$ levels are split by the gyroscopic rotation frequency $\omega$. Note that the resonance frequency is also Doppler shifted by the particle rotation rate since the microwave drive is applied in the lab frame rather than in the rotating frame of the NV center [55, 56, 57].
(b) Barnett anti-alignment— The described setup can be used to observe that the Barnett field $-\omega/\gamma_{0}$ tends to anti-align prolate particles with their own rotation axis [11]. This follows from considering the hard magnet limit $|\omega-\gamma_{0}B|\ll D_{\rm nv}$ where the spin is rigidly attached to the body frame $S_{1}=\hbar m$ with $m\in\{-1,0,1\}$, and $S_{2}=S_{3}=0$. Then the adiabatic Hamiltonian (Gyroscopically stabilized quantum spin rotors) implies that the particle revolves according to the spin-dependent potential $V_{m}(\gamma)=\hbar m(\omega-\gamma_{0}B)\cos\gamma$ [see Fig. 1(c)],
@@
For $B=0$ and $m=1$, the rotor tends to orient with $\gamma=0$ so that the rotation axis anti-aligns with the spin axis. A finite magnetic field $B>0$ weakens the impact of the Barnett field and exactly compensates it at $B=\omega/\gamma_{0}$, where the particle evolves freely in $\gamma$, as would be expected for an object carrying no spin degrees of freedom. Experimentally observing the anti-alignment of the particle spin axis with the rotation axis is thus a unique signature of strong spin-rotation coupling in a rotating nanoparticle. It can be observed in state-of-the-art experiments with nanodiamonds spinning in quadrupole ion traps. This is illustrated in Fig. 2(a), which shows the steady-state alignment $\langle\cos\gamma\rangle$ of the NV axis as a function of the rotation rate $\omega$ for three different temperatures.
(c) Stabilization of spin superpositions— For moderate magnetic fields, the superposition of the magnetic spin states $|\sigma_{x}=+1\rangle=(|S_{1}=+\hbar\rangle+|S_{1}=-\hbar\rangle)/\sqrt{2}$ becomes stabilized when the spin quantization axis lies in the rotation plane. Specifically, for $|\omega-\gamma_{0}B|\lesssim D_{\rm nv}$, the non-magnetic state $|S_{1}=0\rangle$ remains far detuned from the magnetic states $|S_{1}=\pm\hbar\rangle$ for all orientations $\gamma$. The former can thus be adiabatically eliminated, yielding the spin-rotor Hamiltonian in the magnetic spin subspace
@@
as depicted in Fig. 2(b). They exhibit avoided crossings at $\gamma=\pi/2$ (and $23\pi/$ 2), in addition to the potential minima at $\gamma=0,\pi$ for $|S_{1}=\pm\hbar\rangle$ already observed in the hard magnet limit. The energies can be linearized near the avoided crossing yielding $\hbar\Omega_{\pm}(\gamma)=\pm\hbar\delta/2\pm\hbar g^{2}(\gamma-\pi/2)^{2}/\delta$, showing that the spin-rotor is harmonically trapped in the spin state $|\sigma_{x}=+1\rangle$. The groundstate of the resulting harmonic potential is stably trapped for $|\delta|\gg|g\sigma_{\gamma}|$, where $\sigma_{\gamma}=(\hbar\delta/8I_{\rm eff}g^{2})^{1/4}$ is the groundstate width in the effective harmonic potential (see App. E). The state becomes unstable if this condition is violated, as illustrated in Fig. 2(c) for three different magnetic field strengths, or when the spin superposition decoheres after the $T_{2}$ time.
(d) Spin-controlled quantum interference— For strong magnetic fields, $|\omega-\gamma_{0}B|\simeq D_{\rm nv}$, the quantum state of the NV spin can be used to generate and read out superpositions of the angle $\gamma$. Specifically, the magnetic state $|S_{1}=+\hbar\rangle$ can be adiabatically eliminated for small librations around $\gamma\simeq 0$, restricting the spin to the two levels $\left|\uparrow\right\rangle=|S_{1}=-\hbar\rangle$ and $\left|\downarrow\right\rangle=|S_{1}=0\rangle$ with detuning $\Delta=D_{\rm nv}-g$. For small amplitudes and detunings $\langle\gamma^{2}\rangle\ll|\Delta|/|g|\ll 1$, the two spins are effectively decoupled and the resulting dispersive spin-rotor Hamiltonian follows as (App. F)
@@
[…]

### v2

@@ Abstract
The free rotation of a magnetized body can differ significantly from that of the same object if it were nonmagnetic. This, perhaps surprising, discrepancy is due to the fact that magnetic moments carry a spin angular momentum that contributes to the total angular momentum of the magnetized body, as quantified by the material-specific gyromagnetic ratio. Remarkably, it is the total angular momentum, i.e. the sum of mechanical and spin angular momentum, that enters the equations of motion of the body [1]. This gives rise to the so-called spin-rotation coupling [2] between rigid-body rotations and magnetic degrees of freedom, as famously witnessed in the Einstein-de Haas and Barnett effects [3, 4, 5, 6, 7]. The impact of spin-rotation coupling typically decreases with particle size, while it is still relevant for microscale and nanoscale magnetic objects [8, 9, 10, 11, 12, 13].
@@
Nanoscale particles of various shapes and with sizes ranging from hundreds of nanometers to tens of microns can be stably suspended in ultrahigh vacuum using optical, electrical, or magnetic traps [14]. Monitoring and controlling their linear and rotational motion with laser fields enables precision sensing of forces and torques as required for fundamental tests of physics [15, 16]. Recent demonstrations of cooling their center-of-mass oscillations [17, 18, 19, 20] and their angular librations [21, 22, 23] in an optical tweezer into the quantum regime constitute the first steps toward quantum applications of levitated particles. The demonstrated degree of control and readout renders these systems ideal for performing quantum experiments with objects of unrivaled masses [24, 25, 26, 27, 28, 29].
@@
Figure 1: (a) Charged spheroidal nanodiamond with an embedded nitrogen-vacancy (NV) center is levitated in a Paul trap and revolves rapidly with frequency $\omega$ in the symmetry plane of the trap. The major principal axis $\mathbf{n}_{1}$ of the body is chosen to be aligned with the NV quantization axis so that the NV spin states $|S_{1}=\pm\hbar\rangle$ are split by the rotation-induced Barnett field as well as by an applied magnetic field $\mathbf{B}$. The particle rotation induces (b) an effective gyroscopic potential $I\omega^{2}/2\sin^{2}\beta$ stabilizing the out-of-plane rotation angle $\beta$, up to small librations $\xi$, while the external magnetic field together with spin-rotation coupling induce (c) a spin state-dependent potential $V_{m}(\gamma)$ for $\gamma$ rotations around the body-fixed symmetry axis depending on the spin states $|S_{1}=\hbar m\rangle$ with $m\in\{-1,0,1\}$.
@@
Experiments with levitated magnetic objects aim at harnessing their magnetic properties to gain better control over the internal and external degrees of freedom of the nanoparticle [30]. Magnetized particles have been trapped in electric [31, 32], magnetic [33, 34, 35], optical [36, 37], and superconducting traps [38, 39, 40], where their center-of-mass and rotational motion can be monitored and manipulated. Various schemes have been devised to control their mechanical motion via driving the internal magnetization field [41, 42, 43, 44, 45, 46] or individual magnetic defect centers, such as nitrogen-vacancy (NV) centers [47, 48, 30, 49, 50, 51, 52, 53, 54]. These schemes promise to enable sensing of magnetic fields and accelerations [55, 9, 39, 56, 57] as well as the creation of massive superpositions [58, 27, 59, 60, 61, 62], and perhaps even probing the quantumness of gravity [63].
@@
In this Letter, we show how the spin-rotation coupling between a levitated nanoparticle and a single embedded spin degree of freedom can be observed and exploited for quantum experiments. The proposed scheme is based on the recent experimental demonstration of rapid spinning of electrically levitated nanodiamonds with embedded spin degrees of freedom [64, 32]. This spinning boosts the spin-rotation coupling while it also gyroscopically stabilizes the nanoparticle orientation in the corotating frame (see Fig. 1), so a combination of magnetostatic fields and microwave pulses then enables controlling the collective motion of hundreds of millions of atoms via a single spin.
@@
Specifically, we consider a symmetric nanodiamond with a single embedded NV spin levitated in the presence of an external static magnetic field $\mathbf{B}=B\mathbf{e}_{z}$, as schematically depicted in Fig. 1. The nanodiamond is approximated as an ellipsoidal rigid rotor with semiaxis lengths $l_{1}=l_{2}<l_{3}$ and moments of inertia $I_{1}=I_{2}=I$ with $I_{3}<I$. The nanoparticle orientation $\Omega=(\alpha,\beta,\gamma)$, parametrized by Euler angles in the $y^{\prime}-z′′z^{\prime\prime}$ convention, specifies the tensor ${\rm R}(\Omega)$ which rotates from the lab-frame axis ${\bf e}_{j}$, $j=x,y,z$, to the rotor principal axes $\mathbf{n}_{\mu}(\Omega)={\rm R}(\Omega){\bf e}_{j}$, $\mu=1,2,3$. For instance, ${\bf n}_{1}(\Omega)={\rm R}(\Omega){\bf e}_{x}$, etc.; an explicit expression for ${\bf n}_{\mu}(\Omega)$ is given in App. A. The inertia tensor depends on the orientation of the particle through $\mathrm{I}(\Omega)=\sum_{\mu}I_{\mu}\mathbf{n}_{\mu}(\Omega)\otimes\mathbf{n}_{\mu}(\Omega)$. A single NV center with spin-one angular momentum vector operator $\mathbf{S}$ is embedded in the nanodiamond, effectively introducing an internal spin triplet and adding the magnetic dipole moment $\gamma_{0}{\bf S}$ to the particle (with gyromagnetic ratio $\gamma_{0}>0$). The spin angular momentum adds to the mechanical angular momentum to yield the total angular momentum $\bf J$} of the system. For the spin quantization axis, we choose $\mathbf{n}_{1}(\Omega)$, effectively breaking the symmetry of the particle shape; the NV zero-field splitting is $D_{\rm nv}\simeq 2\pi\times 2.87$ GHz. Possible imperfections, such as weak deviations from axial symmetry or small misalignment of the NV axis relative to the principal inertia axis, are analyzed in App. D.
@@
Gyroscopic stabilization is achieved by trapping a charged particle in an azimuthally symmetric quadrupole trap and spinning it with high rotation frequency $\omega$ around the lab-frame ${\bf e}_{z}$ axis, as recently realized in two experiments [64, 32]. For homogeneously charged particles the electric quadrupole tensor aligns with the inertia tensor, with quadrupole moments $Q_{1}=Q_{2}=Q$ and $Q_{3}>Q$. For vanishing electric dipole moment, the center-of-mass and rotational motion of the particle decouple so that the resulting rotational macromotion potential is determined by the ac voltage $U_{\rm ac}$, the electrode separation $d_{0}$, and the ac frequency $\omega_{\rm ac}$ [65]. The spin-rotor Hamiltonian is
@@
Here, $J_{\mu}={\bf J}\cdot{\bf n}_{\mu}(\Omega)$ denotes the body-frame component of the total angular momentum vector and likewise for $\bf S$}. The confining quadrupole potential depends only on the polar angle $\beta$, implying that the lab-frame angular momentum component $p_{\alpha}={\bf J}\cdot{\bf e}_{z}$ is conserved. It is thus convenient to transform the spin-rotor state $|\psi\rangle$ to a corotating frame of boosted angular momentum via
@@
This adds the gyroscopic potential $I\omega^{2}/2\sin^{2}\beta$ [see Fig. 1(b)] to the quadrupole potential in Eq. (Gyroscopically stabilized quantum spin rotors), effectively confining the particle to the rotation plane $\beta\simeq\pi/2$ with harmonic frequency
@@
For large rotation rates $\omega_{\xi}\simeq\omega$ (see App. B), librations out of the rotation plane can be first linearized in $\xi=\pi/2-\beta$ and then adiabatically eliminated. Specifically, linearization yields
@@
with $p_{\gamma}=J_{3}$, $H_{\xi}=p_{\xi}^{2}/2I+I\omega_{\xi}^{2}\xi^{2}/2$, and nondominant contributions to spin-rotation coupling as well as dispersion in the gyroscopic motion being neglected (see App. B). The fast librations out of the rotation plane can now be eliminated by neglecting transitions between eigenstates $|n\rangle$ of $H_{\xi}$ in a Born-Oppenheimer approximation (see App. C). This is achieved by expanding
@@
Figure 2: (a) Steady-state alignment $\langle\cos\gamma\rangle$ as a function of the applied magnetic field. The variance is indicated by the shaded region. (b) Potential energy surfaces $\Omega_{+}$ (solid) and $\Omega_{-}$ (dashed) for $B=-0.5\,\mathrm{mT}$ and $\omega/2\pi=1\,\mathrm{MHz}$. The inset shows the avoided crossing around $\gamma=\pi/2$ where the spin-rotor state $|\chi_{+}\rangle$ can be stably trapped in the $|\sigma_{x}=+1\rangle$ spin state. (c) Spin transition probability for an initial state in spin $|\sigma_{x}=+1\rangle$ (solid) and spin $|\sigma_{x}=-1\rangle$ (dashed) as a function of $\omega_{\gamma}t$ for different magnetic field strengths. The nanodiamond has an ellipsoidal shape with semiaxes $l_{3}=200$ nm and $l_{1}=l_{2}=0.3l_{3}$.
@@
$+\frac{1}{2}(\omega-\gamma_{0}B)\left[f_{n}(S_{1}-iS_{2})e^{-i\gamma}+{\rm h.c.}\right]$
@@
This Hamiltonian describes the coupling between the NV spin and $\gamma$ rotations with effective inertia $I_{\rm eff}=II_{3}/(I-I_{3})$. The final two terms are the adiabatic potential (or potential energy surface) for fixed libration state $|n\rangle$, containing the magnetic potential as well as spin-rotation coupling and the Barnett effect on the NV zero-field splitting. The adiabatic terms are weighted by the overlap integrals
$f_{n}=$ $\langle n|\exp\left(-\frac{ip_{\xi}}{\hbar}\frac{\hbar(\omega-\gamma_{0}B)}{I\omega^{2}}\right)|n\rangle,$ (7a)
$g_{n}=$ $\langle n|\exp\left(-\frac{ip_{\xi}}{\hbar}\frac{2\hbar\gamma_{0}B}{I\omega^{2}}\right)|n\rangle,$ (7b)
for transitions between displaced and nondisplaced libration states. Note that $f_{n}\simeq 1$ for $n\ll I^{2}\omega^{4}/\hbar^{2}\gamma_{0}^{2}B^{2}$ and likewise $g_{n}\simeq 1$. Given that $\hbar\gamma_{0}B/I\omega^{2}\ll 1$ for realistic magnetic fields and rotation rates, the Hamiltonian (Gyroscopically stabilized quantum spin rotors) becomes approximately independent of $n$. Thus, all adiabatic spin-rotor states $|\chi_{n}\rangle$ evolve according to the same effective Hamiltonian:
@@
This Hamiltonian is the main result of this work and the starting point for the proposals discussed below. It accounts for spin-rotation coupling in three different ways. First, the planar $\gamma$ rotation occurs with the effective moment of inertia $I_{\rm eff}$ because $\xi$ librations are gyroscopically stabilized and adiabatically eliminated. Second, the rotation rate $\omega$ of the particle induces the effective spin-dependent potential in the second line of the Hamiltonian (Gyroscopically stabilized quantum spin rotors). Third, the $\mathbf{B}$ field induces a coupling of the spin to rotations around the symmetry axis of the particle as described by the last term in the first line. Note that this coupling is typically small and will be neglected in the following (see App. C). Next, we will study the implications of gyroscopic stabilization for the experimental observation of strong spin-rotation coupling.
(a) Barnett splitting— The Hamiltonian shows how the degeneracy of the $|S_{1}=\pm\hbar\rangle$ states is lifted by the Barnett effect for $B=0$. Specifically, neglecting $\gamma$ rotations by setting $\gamma=0$ and $p_{\gamma}=0$, the $|S_{1}=\pm\hbar\rangle$ levels are split by the gyroscopic rotation frequency $\omega$. Note that the resonance frequency is also Doppler shifted by the particle rotation rate since the microwave drive is applied in the lab frame rather than in the rotating frame of the NV center [66, 67, 68].
(b) Barnett alignment— The described setup can be used to observe that the Barnett field $\omega/\gamma_{0}$ tends to align prolate particles with their own rotation axis [11]. This follows from considering the hard magnet limit $|\omega-\gamma_{0}B|\ll D_{\rm nv}$, where the spin is rigidly attached to the body frame $S_{1}=\hbar m$ with $m\in\{-1,0,1\}$ and $S_{2}=S_{3}=0$. Then the adiabatic Hamiltonian (Gyroscopically stabilized quantum spin rotors) implies that the particle revolves according to the spin-dependent potential $V_{m}(\gamma)=\hbar m(\omega-\gamma_{0}B)\cos\gamma$ [see Fig. 1(c)],
@@
For $B=0$ and $m=1$, the rotor tends to orient with $\gamma=\pi$ so that the rotation axis aligns with the spin axis. A finite magnetic field $B>0$ weakens the impact of the Barnett field and exactly compensates it at $B=\omega/\gamma_{0}$, where the particle evolves freely in $\gamma$, as would be expected for an object carrying no spin degrees of freedom. Experimentally observing the alignment of the particle spin axis with the rotation axis is thus a unique signature of strong spin-rotation coupling in a rotating nanoparticle. It can be observed in state-of-the-art experiments with nanodiamonds spinning in quadrupole ion traps. This is illustrated in Fig. 2(a), which shows the steady-state alignment $\langle\cos\gamma\rangle$ of the NV axis as a function of the rotation rate $\omega$ for three different temperatures. Note that a robust implementation in the presence of rotor asymmetries uses nanodiamonds hosting multiple NV centers (see App. D).
(c) Stabilization of spin superpositions— For moderate magnetic fields, the superposition of the magnetic spin states $|\sigma_{x}=+1\rangle=(|S_{1}=+\hbar\rangle+|S_{1}=-\hbar\rangle)/\sqrt{2}$ becomes stabilized when the spin quantization axis lies in the rotation plane. Specifically, for $|\omega-\gamma_{0}B|\lesssim D_{\rm nv}$, the nonmagnetic state $|S_{1}=0\rangle$ remains far detuned from the magnetic states $|S_{1}=\pm\hbar\rangle$ for all orientations $\gamma$. The former can thus be adiabatically eliminated, yielding the spin-rotor Hamiltonian in the magnetic spin subspace
@@
as depicted in Fig. 2(b). They exhibit an avoided crossing at $\gamma=\pi/2$ (and $3\pi/2$), in addition to the potential minima at $\gamma=0,\pi$ for $|S_{1}=\pm\hbar\rangle$ already observed in the hard magnet limit. The energies can be linearized near the avoided crossing yielding $\hbar\Omega_{\pm}(\gamma)=\pm\hbar\delta/2\pm\hbar g^{2}(\gamma-\pi/2)^{2}/\delta$, showing that the spin rotor is harmonically trapped in the spin state $|\sigma_{x}=+1\rangle$. The ground state of the resulting harmonic potential is stably trapped for $|\delta|\gg|g\sigma_{\gamma}|$, where $\sigma_{\gamma}=(\hbar\delta/8I_{\rm eff}g^{2})^{1/4}$ is the ground state width in the effective harmonic potential (see App. F). The state becomes unstable if this condition is violated, as illustrated in Fig. 2(c) for three different magnetic field strengths, or when the spin superposition decoheres.
(d) Spin-controlled quantum interference— For strong magnetic fields, $|\omega-\gamma_{0}B|\simeq D_{\rm nv}$, the quantum state of the NV spin can be used to generate and read out superpositions of the angle $\gamma$. Specifically, the magnetic state $|S_{1}=+\hbar\rangle$ can be adiabatically eliminated for small librations around $\gamma\simeq 0$, restricting the spin to the two levels $\left|\uparrow\right\rangle=|S_{1}=$-$\hbar\rangle$ and $\left|\downarrow\right\rangle=|S_{1}=0\rangle$ with detuning $\Delta=D_{\rm nv}-g$. For small amplitudes and detunings $\langle\gamma^{2}\rangle\ll|\Delta|/|g|\ll 1$, the two spins are effectively decoupled and the resulting dispersive spin-rotor Hamiltonian follows as (App. G)
@@
Figure 3: (a) Pulse sequence and state evolution of the interference protocol. (b) Probability to measure $\left|\uparrow\right$> at the end of the interferometer sequence as a function of interferometer time $\tau\omega_{\gamma}/\pi$ (evaluated at $\omega/2\pi=1\,\mathrm{MHz}$) for $B=-95\,\mathrm{mT}$ and different rotation frequencies $\omega/2\pi=1\,\mathrm{MHz}$ (gray, solid), $50\,\mathrm{MHz}$ (light blue, dotted), $100\,\mathrm{MHz}$ (dark blue, dashed). The inset shows the probability close to rephasing. (c) Duration of the full protocol (dark blue, solid) and recurrence probability (light blue, dotted) depending on $g/D_{\mathrm{nv}}$. The shaded region indicates when the interferometer time is exceeded by the $T_{2}$ time of the NV spin. The nanodiamond has an ellipsoidal shape with semiaxes $l_{3}=100$ nm and $l_{1}=l_{2}=0.2l_{3}$.